OpenStudy (anonymous):

Verify the identity. 1 + sec^2xsin^2x = sec^2x can someone help me

OpenStudy (bibby):

simplify sec^2x into something using cosine recall that sin/cos = tan and that cos/sin = cot

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

hey super

OpenStudy (anonymous):

i could use the help

Nnesha (nnesha):

do you know trig identities ?? sec = ?? sin = ?? cos = ??

OpenStudy (anonymous):

no not exactly

OpenStudy (bibby):

learn your damn identities you don't want help

OpenStudy (anonymous):

i need someone to walk me through this listen the truth is im taking the class but im running out of time and need months to learn pre calc but only have a week before i cant do school work anymore sorry but i need this done so im open to patially doin this but i need someone to help me with it too

Nnesha (nnesha):

hmm should be on ur notes well \[\large \rm sec = \frac{ 1 }{ \cos }\] \[\csc = \frac{ 1 }{ \sin }\] \[\cot = \frac{ 1 }{ \tan}\] reciprocal functions: \[\cos = \frac{ 1 }{ \sec}~~~~\] \[\sin = \frac{ 1 }{ \csc}\] \[\tan = \frac{ \sin }{ \cos }\] memorize it write these "n ur notebook

Nnesha (nnesha):

so you have to prove that RHS = RHS \[\huge\rm 1+\sec^2x~sin^2x\] change sec^2 by its reciprocal

OpenStudy (anonymous):

@Nnesha ok i copied those functions in my notes

Nnesha (nnesha):

good :)

Nnesha (nnesha):

okay look at the identities find sec^2 = what ??

OpenStudy (anonymous):

sec^2=1^2/cos^2?

Nnesha (nnesha):

yes right 1^2 =1 so we can write \[\sec^2 = \frac{ 1 }{ \cos^2}\]

OpenStudy (anonymous):

so would sec^2(x)-1/cos^2(x)

Nnesha (nnesha):

\[\huge\rm 1+\sec^2x~sin^2x\] replace sec^2 by 1/cos^2

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @quick24 so would sec^2(x)-1/cos^2(x) \(\color{blue}{\text{End of Quote}}\) if you meant "equal" sign then yes

OpenStudy (anonymous):

yes thats what i meant so my equation becomes 1+cos^2(x)sin^2(x)=1/cos^2(x)

Nnesha (nnesha):

well not really bec we are solving LHS to prove that is equal to RHS

OpenStudy (anonymous):

so we leave the right hand side the same

Nnesha (nnesha):

like at the right side there is only one term which is sec^2 yes right stays same bec we havve to prove that LHS = RHS (which is sec^2x)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so my left hand side should like like this 1+cos^2(x) 1/csc^2(x)

OpenStudy (anonymous):

sorry 1+1/cos^2(x) 1/csc^2(x)

Nnesha (nnesha):

nope